### v2

# Resolving thermal gradients and solidification velocities during laser melting of a refractory alloy

Abstract
Metal additive manufacturing (AM) processes, such as laser powder bed fusion (L-PBF), can yield high-value parts with unique geometries and features, substantially reducing costs and enhancing performance. However, the material properties from L-PBF processes are highly sensitive to the laser processing conditions and the resulting dynamic temperature fields around the melt pool. In this study, we develop a methodology to measure thermal gradients, cooling rates, and solidification velocities during solidification of refractory alloy C103 using in situ high-speed infrared (IR) imaging with a high frame rate of approximately 15,000 frames per second (fps). Radiation intensity maps are converted to temperature maps by integrating thermal radiation over the wavelength range of the camera detector while also considering signal attenuation caused by optical parts. Using a simple method that assigns the liquidus temperature to the melt pool boundary identified ex situ, a scaling relationship between temperature and the IR signal was obtained. The spatial temperature gradients ($dT/dx$), heating/cooling rates ($dT/dt$), and solidification velocities ($R$) are resolved with sufficient temporal resolution under various laser processing conditions, and the resulting microstructures are analyzed, revealing epitaxial growth and nucleated grain growth. Thermal data shows that a decreasing temperature gradient and increasing solidification velocity from the edge to the center of the melt pool can induce a transition from epitaxial to equiaxed grain morphology, consistent with the previously reported columnar to equiaxed transition (CET) trend. The methodology presented can reduce the uncertainty and variability in AM and guide microstructure control during AM of metallic alloys.
keywords:
Metal additive manufacturing , High-speed infrared thermography , Infrared to temperature conversion , Grain morphology, Temperature gradient, Solidification velocity, Refractory alloy
PACS:
$0$ , $1$
MSC:
$0$ , $1$
journal: Additive Manufacturing
[…]
[inst1]organization=Department of Mechanical Engineering, University of California, Santa Barbara,postcode=93106, state=CA, country=USA
[…]
[inst2]organization= Materials Department, University of California, Santa Barbara,postcode=93106, state=CA, country=USA
{graphicalabstract}
{highlights}
We present a methodology to convert high-speed IR images during the laser melting of an Nb-base alloy to temperature maps.
We present a method to resolve thermal parameters with high temporal and spatial resolution.
Our thermal data validate the transition from epitaxial to equiaxed grain morphology in a refractory alloy.
1 Indroduction
Laser powder bed fusion (L-PBF), also known as selective laser melting (SLM), is a widely utilized metal additive manufacturing (AM) technique that employs a laser beam to melt fine powders onto a substrate surface, enabling layer-wise and location-specific deposition [1]. This method enables the production of intricate geometries and unconventional materials unattainable through conventional manufacturing, with site-specific control of material properties through tailored printing parameters [2, 3]. The density and properties of the printed material are highly sensitive to the solidification conditions, including the dynamic temperature field. During L-PBF, the high-energy focused laser beam causes rapid melting and high thermal gradients. The resulting heat and fluid flow affect the melt pool dynamics and the cooling rate, which directly control grain growth and resulting microstructures [4]. Unfavorable thermal conditions can lead to high crack susceptibility [5, 6], lack of fusion [7, 8], distortion with surface roughness [5, 8], hot tearing [9], balling [10, 11], and keyholing [12, 11, 13, 14], which can all impact the final density and mechanical properties of the printed component. Previous efforts to optimize thermal processes in L-PBF have mostly relied on trial-and-error prints with varying laser parameters such as power and scan rate [15, 16, 17]. However, only tuning laser parameters can still yield dramatically different as-printed densities, attributed to machine-to-machine variability. Although computational simulations are also used to estimate thermal gradients and solidification velocities during laser melting, these methods are often computationally expensive and can be inaccurate due to the complex multiphysics and multiscale phenomena involved in SLM as well as the lack of thermophysical material properties [18]. As such, microstructure control strategies have relied on variation of process parameters, coupled with computational simulations to estimate the thermal gradients and solidification velocities [19, 20]. Direct measurements of the temperature gradient, cooling rate, and solidification velocity during L-PBF are essential for repeatable microstructure control and defect mitigation [21, 22, 23]. A framework for measuring these values will accelerate development of strategies for microstructure control in new materials during L-PBF.
Thermographic measurements in L-PBF offer valuable insights into the solidification conditions during the manufacturing process. These measurements aid in identifying defect signatures, such as keyholing [14], microstructure prediction [21], and even recognizing dynamic phenomena such as vapor plume dynamics, spatter formation, and thermal history [24, 25, 22, 23]. The complexity of thermal-fluid models also highlights the critical need for precise temperature measurements within and around the melt pool for model validation [26, 6]. For this, various thermographic methods have been utilized, including pyrometers [27, 28], infrared (IR) imaging [25, 25, 29, 22, 23], and two-color method with a color camera [30, 31]. While pyrometry usually provides a spatially averaged temperature, the two-color method is accompanied by a relatively high uncertainty ($\pm$ 100 K). IR imaging has the potential to provide accurate temperature maps, but, it requires calibration of emissivity around the melting temperature, which is particularly challenging for high-temperature materials like refractory alloys [21, 23]. Thus, IR techniques have mainly been used to either qualitatively capture hot spots, melt pool shapes and defects, or estimate temperature using emissivity values calibrated at a lower temperature. More recently, high-speed X-ray imaging has been used synchronously with IR imaging to assign the liquidus temperature to the melt pool boundary identified by X-ray [32, 12, 13, 23, 14]. The emissivity thus characterized was assumed constant around the melt pool to obtain temperature maps based on the Stefan-Boltzmann Law. Despite the advancement, such methods require sophisticated instrumentation in addition to approximating the temperature field because IR cameras do not capture radiation in the full wavelength spectrum as applied in the Stefan-Boltzmann Law. Consequently, there is still a need for a simple and accurate temperature measurement methodology.
Additionally, many previous works have employed insufficient measurement frame rates, typically on the order of tens to thousands of frames per second, which are not fast enough to resolve cooling rates exceeding $10^{6}-10^{7}$ K/s [33, 34, 23]. The limited frame rates can also hinder the accurate measurement of melt pool width and increase measurement uncertainty [25, 29, 35, 21], which underscore the necessity of high-frame-rate measurements. Thus, high-speed IR imaging ($>$10,000 fps) is crucial for accurately resolving fast-moving thermal gradients and solidification processes, ensuring precise characterization of the thermal history [22, 23].
[…]
Here, we report a methodology of converting in situ IR imaging to temperature maps near the melt pool of refractory alloy C103 without the use of X-ray imaging and analyze how the temporally resolved (15,136 fps) thermal characteristics during solidification affect grain size and morphology (Figure 1). A high-speed IR camera was used to capture the thermal radiation intensity maps near the melt pool during laser melting of C103 at different laser energies and scan speeds. The width of the melt pool was measured ex situ using scanning electron microscopy (SEM), and was used to identify the width of the melt pool boundary in the IR images. The liquidus temperature was then assigned to the melt pool boundary in the width direction in IR images. This facilitates the conversion of IR maps to temperature maps assuming a constant emissivity around the melt pool and by integrating spectral radiation only across the wavelength range detected by the camera, as opposed to the Stefan-Boltzmann scaling applied in previous literature. Using this methodology, the melt pool shape is identified and its moving velocity is analyzed. Thermal characteristics of the solidification process including temperature gradients ($G$), cooling rates ($dT/dt$), and solidification velocities ($R$) were analyzed. Ex situ SEM images reveal two distinct solidification modes, including epitaxial growth through the melt pool boundary and nucleated equiaxed grains within the melt pool. These microstructures are analyzed with respect to local solidification velocities and thermal gradients. This work demonstrates a methodology to resolve detailed time-dependent thermal characteristics during solidification and provides insights into the link between thermal process conditions and solidification behavior during metal AM.
2 Methods
2.1 Single-track laser scanning: experimental setup and procedure
Figure 1: Overview of the experimental procedure. A high-speed IR camera records radiation intensity maps which are then converted to the dynamic temperature maps during laser melting of C103. The temperature maps offer rich spatially and temporally resolved information such as the melt pool shape, the temperature gradient, the heating/cooling rate, and the solidification velocity. The microstructures of the melt pools are analyzed using SEM, indicating different grain growth dynamics under various laser processing parameters. Processing-microstructure relationships can be established by comparing thermal characteristics with the resulting microstructures.
Single-track laser experiments were performed in a custom inert laser processing chamber interfaced with a high-speed IR camera (see Supporting Information S1). The chamber offers the basic functionality of a conventional L-PBF 3D printer for conducting rapid solidification experiments and was modified from an experimental setup in previous work by Mullin et al.[39]. The employed heat source is an IPG Photonics 400 W Nd:YAG continuous laser operating at a wavelength of 1070 nm. Optical components, including a collimator, mirror galvanometer, and focus lens, are integrated to collectively manipulate and focus the laser beam. A customized Python code automates single-track laser scanning experiments by controlling both the galvanometers and the laser. The beam is focused down to $\approx$ 80 $\mu$m on the build plate, which is situated on a motorized linear stage in the chamber.
[…]
2.2 Temperature conversion
Figure 2: Conversion of measured intensities (NUC) from the IR camera to temperature. (a) An example of a raw IR image of the melt pool of C103 showing measured IR intensity (NUC) in each pixel (30 ×$\times 34\mu$m) under the laser scanning condition of 170 W and 600 mm/s. The same raw IR intensity map is replotted in (b), where the vertical axis is the magnitude of the NUC value and a smooth function is applied to the curved surface to reduce the coarsening effect due to the low spatial resolution of the IR camera. A horizontal plane is superimposed on this 2D map to identify the melt pool boundary. (c) The melt pool shape extracted from the intersection between the horizontal plane and the intensity map in (b), where the width ($W$) of the melt pool agrees with the width ($W$) shown in ex situ SEM images (inset). (d) The temperature map around the melt pool converted from the raw IR image.
Obtaining accurate temperature maps from IR measurements requires careful calibration because the detected IR signal depends on the spectral emissivity of the sample and signal attenuation through optical windows and mirrors. These are challenging and time consuming to measure at high temperatures, particularly near the liquidus of high temperature alloys like C103 ($T_{\text{liquidus}}\!\approx$ 2411 C). Therefore, recent works have assigned the liquidus temperature to the melt pool boundary identified through synchronous X-ray imaging [22]. The temperatures of other pixels in the image were scaled based on the Stefan-Boltzmann law ($E\!\sim\!T^{4}$) which integrates thermal radiation across the entire spectrum with a gray body assumption. However, since almost all IR cameras only detect radiation within a narrow wavelength band, the Stephan-Boltzmann approach will lead to an inaccurate scaling relationship. In this work, the melt pool boundary is first identified without X-ray imaging by simply identifying the width of the melt pool through ex situ SEM imaging. More importantly, we integrate thermal radiation only within the spectral range of the IR camera (Telops M3K) sensor (3-5.5  $\mu$m), when scaling the temperature map from the liquidus temperature. Figure 2(a) shows the measured raw IR intensities (NUC value) for a laser scanning condition of 170 W and 600 mm/s at a specific time ($t\!=\,$3.6 ms) where $x$-direction is the laser scanning direction (from left to right). The same raw IR intensity map is replotted as shown in Figure 2(b) where the vertical axis is the magnitude of the NUC value and a smooth function (spline in MATLAB) is applied to the curved surface to reduce the coarsening effect due to the low spatial resolution of the IR camera. Since the front half of the melt pool boundary should technically be around the liquidus temperature assuming very small superheating during melting, an arbitrary horizontal plane intersects this curved surface to form an imaginary ”melt pool boundary” where NUC values on the melt pool boundary are all the same representing the same temperature. The vertical position of the horizontal plane is then adjusted until the resulting width of the imaginary melt pool in the $y$-direction ($W$ in Figure 2(c)) agrees with the actual width of the melt pool measured through ex situ SEM imaging (inset of Figure 2(c)). The NUC value or the raw intensity value thus identified (i.e., the final vertical position of the horizontal plane) represents the signal detected by the IR camera corresponding to the liquidus temperature ($T_{\text{liquidus}}\!\approx$ 2411 C) of the sample [39], which already accounts for attenuation through optical parts in the experimental setup. In the example shown in Figure 2(a-c), this critical IR intensity value is identified to be 19720 which is the critical NUC value under the selected camera acquisition settings (20 $\mu$s exposure time, and 15,136 fps frame rate). Here, the liquidus temperature ($T_{\text{liquidus}}$) is used to correlate with the melt pool boundary intensity instead of the solidus temperature because the width of the laser track is determined by the melting process.
To further obtain the temperature map, we first numerically integrate Planck’s law, which describes blackbody radiation, over the 3–5.5 $\mu$m wavelength range, to derive the theoretical emitted energy within this band from the sample surface (Supporting Information Figure S3(a) blue line). The In-Band Irradiance (IBI), which is the blackbody radiation striking the camera detector’s surface with a unit of W/m2, is only a portion of the emitted blackbody radiation due to a view factor (Supporting Information Figure S3(a) red line). The view factor is dependent on the camera viewing angle and distance to the sample which is determined to be 1/8 for our experimental condition (Supporting Information Figure S3(b)). A 9th-order polynomial function is fitted to describe the numerical relationship between the IBI and temperature (Supporting Information Figure S3(b)). This relationship roughly describes how the detected IR raw signal scales with temperature for a blackbody, which differs from the $E\!\sim\!T^{4}$ relationship in the Stefan-Boltzmann Law. The non-unity in-band emissivity of the sample, non-ideal spectral absorbance of the sensor, and transmittance of the optics in the IR camera require further correction to convert the theoretical IR signal (IBI) to the actual detected IR signal (NUC). To address this, the camera software (Reveal IR) offers a built-in correction to convert the IBI to NUC, which is also influenced by the camera acquisition setting, such as exposure time and frame rate. Specifically, for our experimental camera setting, as explained in Section 2.1, the relationship between NUC and IBI is introduced below.
$\text{NUC}=\varepsilon\times 21.045\times\text{IBI}+1371.2$ (1)
where $\varepsilon$ is the effective emissivity of the C103 alloy, which accounts for the intrinsic in-band emissivity of C103 and additional signal attenuation from the gold mirror and the Ge window. The final calibration curve between the NUC and temperature, obtained using this conversion method, is shown in Figure S3(c) (Supporting Information, Section S3). Based on this NUC vs Tempearure relationship, the effective emissivity of C103 around the melt pool boundary was determined by adjusting its value until the NUC equals the critical NUC value (19720 in Section 2.2) when the temperature is the liquidus temperature of C103 metal. The effective emissivity of C103 at the liquidus temperature thus determined is $\varepsilon$ = 0.083. The total hemispherical emissivity averaged over all wavelengths and hemispherical directions of C103 reported in the literature is 0.28 at 800 C and 0.4 at 1200 C [43]. While the emissivity of C103 in its liquid state is not reported in previous studies, it is generally observed that the emissivity of metals in their liquid state can be lower than in their solid state [44]. In addition, emissivity values are sensitive to surface conditions and smooth and non-oxidized surfaces have lower emissivity than rough and oxidized counterparts. Given that the effective emissivity reported in this work is only the in-band emissivity between 3–5.5 $\mu$m in the camera viewing direction at the liquid-solid transition temperature for a polished C103, and further accounts for attenuation via the gold mirror and Ge window, this value of 0.083 is reasonable.
As a result, this conversion process yields the temperature map of the entire domain as illustrated in Figure 2(d). To obtain a more accurate map, the temperature-dependent emissivity of the molten liquid metal and the solid C103 needs to be measured, which is beyond the scope of this work. However, in this work, since we are mainly interested in characterizing the temperature gradient and the cooling rate at the melt pool boundary, the assumption of a constant emissivity ($\varepsilon\!\approx$0.083) in calculating the temperature near the melt pool boundary is appropriate.
3 Results
3.1 Ex situ microstructure characterization using SEM
Figure 3: Ex situ characterization of the microstructure and dimensions of laser melted C103. (a) SEM images of the top-down view of the printed material after laser scanning. (b) SEM images of the cross-section of the melt pool. (c) The measured width of the melt pool corresponds to the width of the melt pool ($W$) in (a) under each laser parameter. The red bar graph shows the nucleated grains within the melt pool ($W\,\gtrapprox$ 90 $\mu$m), while the blue bar graph shows the epitaxial growth through the melt pool ($W\,\lessapprox$ 90 $\mu$m). (d) The aspect ratios ($W/D$) of the melt pool are compared with the energy density ($\rho_{\text{E}}$) of the input laser. Higher aspect ratio values ($\gtrsim$ 7.8) indicate nucleated grains, while lower aspect ratios ($\lesssim$ 7) indicate epitaxial growth.
Figure 3 shows micrographs of a single-track experiment conducted on a wrought C103 substrate. After scanning the C103 alloy under different laser conditions ($\rho_{\text{E}}\approx$ 0.2 – 0.8 J/mm, where $\rho_{\text{E}}$, the energy density, is the ratio between the laser power and the laser scan speed), the top-down view (Figure 3(a)) and the cross-sectional view (Figure 3(b)) of the samples are imaged using a scanning electron microscope (ThermoFisher, Apreo C) with a backscatter (BSE) detector. The measured widths ($W$) of the scanned melt pool extracted from Figure 3(a) are plotted in Figure 3(c). The cross-sectional view of the laser melted C103 in Figure 3(b) shows the width ($W$), the depth ($D$), and the microstructure of the melt pool. The more detailed SEM images can be found in Supporting Information S4. An intriguing observation is that track melts with a wider melt pool width ($\gtrsim$ 90 $\mu$m) shows nucleated grains within the melt pool, as indicated by the red circle in Figure 3(b). In contrast, narrower melt pool width ($\lesssim$ 90 $\mu$m) resulted in only epitaxial growth from the melt pool boundary, under the few experimental conditions tested. Figure 3(d) illustrates the aspect ratio of the melt pool ($W/D$) alongside the 2D energy density ($\rho_{\text{E}}$) of the scanning laser for the six laser parameters examined. Relatively lower aspect ratio values ($W/D$$\lesssim$ 7) indicate nucleated grains, while relatively higher aspect ratios ($W/D$$\gtrsim$ 7.8) indicate epitaxial growth where the transition is at an aspect ratio of approximately 7–8. Despite the observed trend between melt pool width, aspect ratio, and grain growth behavior, the thermal boundary conditions at which microstructural changes occur have yet to be determined for C103 and, more broadly, for refractory alloys. This is primarily due to the challenges associated with characterizing these values at high melting points in an inert environment. A detailed characterization of spatial and temporal temperature gradients, cooling rates, and solidification velocities during the experiment may shed insights into a more universal and machine-agnostic understanding of processing-microstructure relationships.
Figure 4: Melt pool boundary and its movement over time. (a) 2D temperature map under a laser scanning parameter of 170 W and 600 mm/s. The horizontal plane is the solidus temperature of the C103 alloy. The inset shows the magnified view of the melt pool. The circular line (red and blue dotted line) that intersects the 2D temperature map and the solidus temperature of C103 is the boundary of the melt pool. (b) the melt pool boundary at multiple time steps under various laser scan speeds.
3.2 The melt pool shape and velocity
The time-dependent temperature maps can be further processed to extract the temperature gradient ($G$), the heating/cooling rate ($dT/dt$), the 2D melt pool shape, and the solidification velocity ($R$) along the melt pool boundary, which are all important thermal parameters potentially affecting grain growth [45, 46, 47, 48]. Due to the relatively coarse pixel size of the IR camera (30 $\mu$m and 34 $\mu$m in $x$ and $y$ directions, respectively), the 2D temperature map is first interpolated to form a smoother map (using spline in MATLAB), as shown in Figure 4(a) where the interpolated temperature map was plotted against $x$ and $y$ coordinates. To locate the melt pool boundary, especially the rear half of the melt pool where solidification occurs, a horizontal plane at the solidus temperature of C103 ($T_{\text{solidus}}\!\approx$ 2141 C) was constructed to intersect the interpolated temperature surface to form a closed-loop line that represents the melt pool boundary, as shown in the inset of Figure 4(a). More strictly speaking, it is more appropriate to use the liquidus temperature ($T_{\text{liquidus}}\!\approx$ 2411 C) to identify the front half (melting, as shown in the red dotted line in the inset of Figure 4(a)) of the melt pool and the width of the melt pool, which was employed in the earlier Section 2.2. Since grain growth may occur during solidification, solidus temperature is used for a more accurate depiction of the rear half of the melt pool shape. Here, we also assumed a small subcooling temperature compared to the thermodynamic solidus temperature. This assumption needs to be further investigated due to the high cooling rates employed in the study. Figure 4(b) shows the melt pool boundary at multiple time steps (0, 0.66, 1.32 ms) under various laser scan speeds (200, 400, 600 mm/s) for a laser power of 170 W. As expected, a lower scan speed (200 mm/s) results in a slightly larger melt pool size compared to a higher scan speed (600 mm/s), presumably due to a higher energy density (for a more detailed melt pool shape comparison, see Supporting Information Section S5.). The velocities at which the melt pools move extracted from Figure 4(b) agree well with the laser scanning velocities. The melt pool shape and width remain relatively constant, suggesting that the system is approximately approaching the steady state (see Supporting Information Section S6).
Figure 5: Temporal thermal characteristics along the laser scanning direction. (a) A schematic illustrates the movement of the melt pool during laser scanning in the $x$-direction, with a star marking a reference point along the laser scanning line. (b) The temperature ($T$), (c) temperature gradient ($G_{x}=dT/dx$), and (d) heating/cooling rate ($dT/dt$) of a reference pixel along the laser scanning line over time are depicted. A horizontal dotted line in the temperature map (b) indicates the solidus temperature of the C103 alloy, while a vertical dotted line in (b),(c), and (d) marks the moment of solidification.
3.3 Thermal characterization: temporal analysis
To characterize the temporal temperature response, the following analysis examines a single point along the laser scanning direction($x$-direction) as illustrated in Figure 5(a). The melt pool moves in the $x$-direction over a reference point denoted with a star. As the laser moves, this reference point melts and subsequently solidifies. The temperature ($T$), the temperature gradient ($G$), and the heating/cooling rate ($dT/dt$) of the reference point for a representative laser parameter (170 W, 600 mm/s) are shown in Figure 5(b-d). The circular data points represent the actual data captured by the infrared camera, while the smooth solid line is interpolated from the smoothed 2D temperature profile using the same method as described in Figure 4a. The temperature response in Figure 5(b) shows rapid heating and cooling within only 0.6–1.2 ms, highlighting the necessity for high-speed characterization which is achievable with our camera frame rate. The peak temperature in the melt pool could be underestimated due to an expected lower emissivity of liquid metal than the solid. In addition, the asymmetry of the temporal response of temperature, or more rapid heating than cooling, agrees well with previous thermal simulations and analytical solutions for a moving heat source on a solid substrate [49, 50]. Solidification occurs when the temperature drops to the solidus temperature (dashed horizontal line) during cooling, corresponding to $t=t_{\textbf{l-s}}$ in Figure 5(b).
The temperature gradient on the $x$$y$ plane is
$G=\sqrt{\left(\frac{dT}{dx}\right)^{2}+\left(\frac{dT}{dy}\right)^{2}},$ (2)
Note that the total temperature gradient involves the contribution of a temperature gradient in the $z$-direction, which is not captured by the IR camera. However, only the $x$-direction temperature gradient $G_{x}$ is needed along the center laser scan path due to symmetry in the $y$ direction,
$G_{x}=\frac{dT}{dx}\approx\frac{\Delta T_{x}}{\Delta x},$ (3)
where $\Delta x$ is the distance between adjacent points in the smoothed temperature map in the $x$-direction (3 $\mu$m after smoothing the temperature map), and $\Delta T_{x}$ is estimated using a linear approximation of the temperature centered difference method (for more details about the gradient derivation, see Supporting Information S7).
The temporal response of the temperature gradient ($G_{x}$) (Figure 5(c)) shows two peaks around 3.4 ms and 3.7 ms, which happen during melting and solidification (by identifying the temperatures at 3.4 ms and 3.7 ms in the temperature versus time curve). On the contrary, the temperature gradient within the melt pool is much lower, despite the highest local laser energy flux. This can result from a higher effective thermal conductivity of the liquid due to convective flow, effectively reducing the temperature gradient within the melt pool. In addition, the temperature gradient peak during melting is higher than the temperature gradient peak during solidification, which agrees with the asymmetry in the heating and cooling process. Figure 5(d) also ($dT/dt$) shows more rapid heating ($\approx$ $12\,\times 10^{6}\,$/s) as compared to cooling ($\approx$ $7\,\times 10^{6}$ K/s). The magnitude of thermal gradient upon cooling is within estimated ranges for L-PBF (104 - 108 K/m) [51]. Though thermal gradients under laser melting conditions have not previously been measured for Nb-base or refractory alloys generally, thermal simulations have produced values within the range of 107 K/m for C103 [42] and 107-108 for pure refractory metals with higher thermal conductivity (Ta, Mo) [52, 53]. It is anticipated that the lower effective thermal conductivity and higher absorptivity of powder feedstock for L-PBF may influence the thermal gradients, which demands further investigation.
3.4 Thermal characterization: on the melt pool boundary
Figure 6: Spatial thermal information along the melt pool boundary. (a) The temperature gradient ($G$), (b) cooling/heating rate ($dT/dt$), and (c) solidification velocity ($R$) for 170 W 600 mm/s laser scanning at the 3.6 ms time frame. The laser is moving in the +$x$-direction. Temperature gradient and solidification velocity do not have a $z$-component contribution, as the spatial analysis is based on 2D temperature measurements. The thermal information along the arc from the outer edge to the end of the melt pool (as illustrated in (a)) is also extracted from the spatial thermal data, such as (d) the thermal gradient ($G$), (e) the cooling/heating rate ($dT/dt$), and (f) the solidification velocity ($R$), which are plotted against the arc length ($s$) in different laser scan conditions.
In addition to investigating the temporal response of one point along the laser scanning line, we also characterize the spatial distribution of temperature gradient ($G$), heating/cooling rate ($dT/dt$), and solidification velocity ($R$) on the melt pool boundary, as shown in Figure 6(a-c). For a laser scanning condition of 170 W and 600 mm/s at a fixed time of 3.6 ms, the melt pool boundary is extracted using the same approach as in Figure 4. The temperature gradient ($G$) and solidification velocity ($R$) on the melt pool boundary do not account for a $z$-direction component, as the spatial analysis is based on 2D temperature measurements. The temperature gradient ($G$) was calculated based on Equation 2, which accounts for the gradients in the $x$ and $y$ directions. Its numerical form is
$G\approx\sqrt{\left(\frac{\Delta T_{x}}{\Delta x}\right)^{2}+\left(\frac{\Delta T_{y}}{\Delta y}\right)^{2}},$ (4)
[…]
The solid-liquid interface velocity, $R$, also known as the solidification velocity, can be calculated based on the ratio between the time derivative of the cooling rate ($dT/dt$) and the temperature gradient ($G$), as shown in previous studies [21]
$R=\frac{\left(\frac{dT}{dt}\right)}{G},$ (5)
Similarly, the negative value of $R$ represents solidification, whereas the positive $R$ represents melting. Mathematically, the solidification velocity derived from this equation represents the solid-liquid interface velocity in the temperature gradient direction. In our work, since $G$ does not include a $z$-direction component from the 2D temperature measurement, the $R$ value shown in Figure 6(c) only represents the projected velocity on the $x$-$y$ plane. In Figure 6(a-b), $G$, $dT/dt$ at the trailing edge of the melt pool along the laser scanning path agrees well with values shown in Figure 5(c-d), confirming that our analysis is consistent.
Additionally, the temporal response of the solidification velocity at the melt pool edge can be obtained by tracking the melt pool boundary from the side edge to the center, perpendicular to the laser scanning direction. The solidification velocity in the $y$-direction as a function of time is presented in Supporting Information Section S8. It can be observed that the $y$-direction solidification velocity also accelerates as the edge moves toward the center. Both the gradient-based solidification velocity (Figure 6(c)) and the y-direction solidification velocity (Section S8) illustrate the increase in solidification velocity from the edge (point a in Figure 5(a)) to the center(point b in Figure 5(a)). The terminal velocity of $R$, which is the velocity at the trailing edge, agrees well with the laser scanning velocity, indicating that the system is approximately at the steady state.
3.5 Microstructural transition during solidification
The spatially resolved thermal features in Figure 6(a-c) allow further analysis of the columnar-to-equiaxed transition (CET), observed in four of the laser scanning conditions. Solidification of a melt pool typically progresses from the melt pool boundary toward the center of the melt pool in three dimensions. However, since we can only measure temperature on the top $x$-$y$ plane with no $z$-component, the following analysis is focused on investigating the thermal features ($G$, $dT/dt$ and $R$) leading to the grain morphology transition observed on the top $x$-$y$ plane, specifically from the side edge (point a in Figure 5(a)) to the trailing edge (point b in Figure 5(a)) of the melt pool. To illustrate this transition, thermal features are extracted along the arc ($s$) of the melt pool boundary from the uppermost edge to the trailing edge, as illustrated from ’a’ to ’b’ in Figure 6(a). The corresponding temperature gradient ($G$), cooling rate ($dT/dt$), and solidification velocity ($R$) are then plotted against $s$ in different laser scanning conditions, as shown in Figure 6(d-f).
Figure 6(d) shows that the temperature gradient decreases from the edge to the center of the melt pool, while the cooling rate (Figure 6(e)) and solidification velocity (Figure 6(f)) increase from the edge to the center of the melt pool for all of the laser conditions. One interesting observation is that the cooling rate (Figure 6(e)) changes with the laser scan speed, and is effectively independent of the laser scan power. This result is consistent with findings in previous studies [54, 55]. This is because faster scan speeds reduce the time the laser spends on each location, limiting the material’s ability to absorb and retain heat, while the laser power primarily affects the size and temperature of the molten pool rather than the cooling rate. In particular, the faster scan speed (600 mm/s) results in a higher cooling rate. For the four laser conditions where a CET was observed (Figure 3(b)), epitaxial or columnar growth is observed at the edge, whereas nucleated or equiaxed grains are observed at the center of the melt pool track (see Figure 7(a)). Correlating the spatially resolved thermal data with the location-dependent grain growth suggests that a higher temperature gradient, lower cooling rate, and lower solidification velocity on the edge of the melt pool more likely lead to epitaxial growth. The transition to equiaxed grains occurs as the temperature gradient decreases and the cooling rate and solidification velocity increase towards the center of the melt pool.
Figure 7: Microstructure changes during the solidification. (a) shows SEM images of the cross-section of the melt pool under a laser scanning parameter of 170 W and 600 mm/s. The early stage of cooling exhibits epitaxial growth starting from the edge, while the later stage shows the formation of nucleated grains at the center of the melt pool. (b) represents the thermal phase map between temperature gradient and solidificiation velocity, illustrating the transition from epitaxial growth to nucleated grains. This transition occurs with a decreasing temperature gradient and increasing solidification velocity, consistent with previously reported CET maps [56, 57].
The observed CET for C103 is consistent with other conventional metals, such as Ti–6Al–4V and IN718 [56, 57], where the CET has been plotted on a temperature gradient ($G$) vs solidification velocity ($R$) maps. In previously reported CET maps between $G$ and $R$, the CET curve was calculated through simulation [57], demonstrating that the transition occurs when the temperature gradient ($G$) decreases and the solidification velocity ($R$) increases during solidification. This understanding aligns with our observations for C103, as shown in Figure 7(b). Specifically, Figure 7(b) plots the temperature gradient ($G$) and solidification velocity ($R$) of all locations along the arc of the melt pool, extracted from Figure 6(a) and (c). Data in the top left region in Figure 7(b) (blue shaded area) are located at the outer edge of the melt pool, and the bottom right region (red shaded area) shows data at the center (trailing edge) of the melt pool. The starting and ending points of the solidification are represented as filled data points, which are associated with epitaxial growth and nucleated grains, respectively. This approach provides experimental data to construct a CET map for C103, which provides guidance for microstructure control efforts during SLM.
4 Conclusion
In summary, in situ IR measurements of the Nb-based alloy C103 during laser melting were conducted to extract temporally resolved temperature fields during solidification. Microstructure development was examined ex situ, revealing epitaxial growth and nucleated grains under different laser conditions. To resolve the detailed thermal characteristics leading to the microstructures, we present a calibration methodology to convert the detected IR intensities to temperature maps by carefully considering in-band irradiance, view angle, non-ideality of optics as well as the emissivity of the sample. The melt pool boundary as well as its moving dynamics are presented. We analyze the transient response of a point along the laser scanning path, highlighting the rapid change of temperature, temperature gradients, and cooling rates within only 1 ms. In addition, spatially resolved temperature gradients, cooling rates, and solidification velocity on the melt pool boundary are presented, while the time-dependent solidification process from the melt pool edge inward is also illustrated. This approach enabled the creation of a thermal phase map that associates microstructural transition during the solidification process to the local temperature gradient and solidification rate. This study underscores the importance of in situ IR measurements for a detailed understanding of solidification conditions and microstructure development during additive manufacturing. Moreover, the ability to accurately measure temperature fields with high temporal resolution is crucial for addressing machine-to-machine variability in AM processes, which can lead to significant differences in microstructural outcomes. By enhancing our understanding of these thermal processes, the IR method offers a valuable tool for improving consistency and control across different AM systems.
Acknowledgment: The thermal measurement and analysis part of this work is supported by the National Science Foundation through the Materials Research Science and Engineering Center (MRSEC) at UC Santa Barbara: NSF DMR–2308708 (Seed). The laser melting experimental setup and materials characterization are based upon work supported by the Department of Energy, National Nuclear Security Administration under Award Number(s) DE-NA0004152. O.A. Wander and K.M. Mullin also acknowledge support from the National Science Foundation Graduate Research Fellowship under Grant No. 2139319. We thank Prof. Patrick Hopkins and Hunter Schonfeld at the University of Virginia for discussions on emissivity measurement. We thank Wes Autran, Alex Côté, Vince Morton, Andrew Niehaus, and Hector Moreno from Telops for detailed discussions and suggestions on the temperature conversion process. We thank Alden Wong and Sachin Suri for their contributions to data analysis and development of experimental hardware. We thank S. Farzad Ahmadi for the comments on the figures. The research reported here made use of the shared facilities of the Materials Research Science and Engineering Center (MRSEC) at UC Santa Barbara: NSF DMR–2308708. The UC Santa Barbara MRSEC is a member of the Materials Research Facilities Network (www.mrfn.org).
Disclaimer: This report was prepared as an account of work sponsored by an agency of the United States Government. Neither the United States Government nor any agency thereof, nor any of their employees, makes any warranty, express or implied, or assumes any legal liability or responsibility for the accuracy, completeness, or usefulness of any information, apparatus, product, or process disclosed, or represents that its use would not infringe privately owned rights. Reference herein to any specific commercial product, process, or service by trade name, trademark, manufacturer, or otherwise does not necessarily constitute or imply its endorsement, recommendation, or favoring by the United States Government or any agency thereof. The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof.
These authors contributed equally to this work. Corresponding authors. E-mail addresses: tresap@ucsb.edu, yangying@ucsb.edu
Supporting Information
1. Single-track laser scanning and infrared measurement setup
Figure S1: A custom laser scanning setup interfaced with a high-speed IR camera. (a) a schematic of a custom laser processing chamber and optical components to interface with an infrared camera. (b) a photographic image of the custom chamber and infrared camera, with an inset detailing the mirror mount inside the side plate of the chamber.
Figure S1 shows the schematic (Figure S1(a)) and photographic image (Figure S1(b)) of the custom Argon laser processing chamber interfaced with a high-speed IR camera. The custom laser processing chamber has been previously developed by Mullin et al. and the details can be found in reference [39]. Dynamic temperature maps during laser scanning were captured using a high-speed IR camera (Telops M3K) equipped with a 1X lens (Telops) with a frame rate of 15,136 fps. The IR camera can capture up to 100,000 fps. However, at this extremely high frame rate, the window size is limited to 64 px×$\times$4 px (1.92 mm×$\times$0.12 mm). To capture the entire melt pool and track its movement, a moderate window size (3.84 mm×$\times$2.18 mm) was selected, with a corresponding frame rate of 15,136 fps which is enough to capture the solidification dynamics. The IR camera is positioned adjacent to the laser scanning chamber. It views the sample through a germanium (Ge) window (transparent between 3-5 $\mu$m) and an IR-reflecting gold mirror, allowing the recording of thermal radiation emitted from the melt pool during laser scanning. The received signal via the IR camera is in the form of radiant intensity, necessitating an accurate conversion of the intensity maps into temperature fields before further thermal analysis.
2. Transmittance spectra of the Ge window
Figure S2: Transmittance spectra of the Ge window. The high-speed IR camera (Telops M3K) has a spectral range of 3–5.5 $\mu$m and the germanium (Ge) window is mostly transparent within this range. The non-unity transmittance of the Ge window especially near 3 $\mu$m is accounted for in the calibration and is reflected by the effective emissivity. The transmission data was provided by the Ge window manufacturer (Sinoptix Optical).
3. Temperature conversion
Figure S3: Calibration curve from raw data to temperature
The theoretical irradiance incident on the camera detector’s surface, referred to as the In-Band Irradiance (IBI) value (measured in W/m2), is calculated from Plank’s law as shown below.
$L_{\lambda}(T)=\frac{2hc^{2}}{\lambda^{5}}\cdot\frac{1}{\exp\left(\frac{hc}{\lambda k_{B}T}\right)-1}$ (6)
[…]
To calculate the in-band irradiance received by the camera from the theoretical spectral radiation, several factors must be considered, including the camera’s tilt angle, view factor, and the camera’s solid angle. The IR camera manufacturer, Telops, provides a calibration to account for these parameters, converting temperature to In-Band Irradiance (IBI). The relationship between temperature and IBI is shown in Figure S3(a) as red data points.
[…]
Factors such as the sample’s non-unity in-band emissivity within 3-5.5 $\mu$m, the sensor’s non-ideal spectral absorbance, and the optics’ transmittance in the IR camera necessitate additional corrections to convert the theoretical IBI to the actual detected signal (NUC). The camera software (Reveal IR) includes a built-in correction to perform this conversion, taking into account camera acquisition parameters like exposure time and frame rate. Details on how the NUC and IBI are related under our experimental condition, are described in the main text equation Equation 1. Figure S3(c) shows the final result of the temperature vs NUC curve.
4. Scanning electron microscope image
Figure S4: SEM images of the cross-section of the melt pool. The wider melt pool width shows nucleated grains within the melt pool (red boxes), in contrast, only epitaxial growth from the melt pool boundary was observed for melt pools with narrower widths (blue boxes).
5. Melt pool shape
Figure S5: The solid-liquid boundary shape of each experiment is determined using the method explained in Figure 4 in the main text. The higher power (170 W) results in a larger melt pool size compared to the lower power laser (130 W) at the same scan speed. Also, the slower speed (200 mm/s) yields a larger melt pool size than the higher speed (600 mm/s) at the same laser power. Due to a hardware issue, the IR recording of the 130 W 600 mm/s was not saved correctly.
6. Isothermal line velocity
Figure S6: Comparison of the melt pool moving velocity measured from Fig. 4(b) with the solidification velocity of the trailing edge of the melt pool calculated using Equation 5 in the main text(black data).
The cooling rate ($dT/ds$) and the temperature gradient ($G$) of a point on the melt pool boundary can be used to calculate the solidification velocity. As explained in the main text, typically the total temperature gradient includes a contribution from the $z$-direction. Since in this paper, $G$ is the projected temperature gradient on the $x$-$y$ plane, the solidification velocity calculated is the projected solidification velocity on the $x$-$y$ plane. For the trailing edge on the melt pool boundary, $R_{x}$ (blue data) is the solidification velocity calculated from Equation 5 in the main manuscript, which represents the isothermal line velocity. The results align well with the isothermal line velocity obtained by directly tracking the movement of the melt pool boundary, as indicated in Figure 4(b) of the main text (black data point).
7. Temperature gradient calculation
Solidification occurs when the temperature drops to the solidus temperature during cooling, assuming minimal subcooling, which needs further verification. The magnitude of the temperature gradient on the $x$$y$ plane is
$G=\sqrt{\left(\frac{dT}{dx}\right)^{2}+\left(\frac{dT}{dy}\right)^{2}},$ (7)
The total temperature gradient in 3D should include a $z$-direction component. However, since the IR camera only captures temperature maps on the $x$-$y$ plane with no $z$-direction information, the temperature gradient we are considering is only the projected gradient on the $x$-$y$ plane. For the temperature gradient along the center laser scan path, only the $x$-direction temperature gradient $G_{x}$ is needed due to symmetry in the $y$-direction, so Equation 7 reduced to
$G_{x}=\frac{dT}{dx}\approx\frac{\Delta T_{x}}{\Delta x},$ (8)
where $\Delta x$ is the distance between adjacent points in the smoothed temperature map in the $x$-direction (3 $\mu$m after smoothing the temperature map), and $\Delta T_{x}$ is estimated using a linear approximation of the temperature centered difference method which is
$\Delta T_{x}=\frac{\mid T_{x}-T_{x-1}\mid+\mid T_{x}-T_{x+1}\mid}{2},$ (9)
In Equation 9, $x$ is the corresponding point location in the $x$-direction of the image, and $T_{x+1}$ and $T_{x-1}$ are the temperatures of adjacent pixels. The discretized temperatures (every 3 $\mu$m) were obtained by interpolating the original temperature data points, which have a coarser mesh, using a spline smooth function in MATLAB.
For the temperature gradient map on the melt pool boundary, the numerical form of equation 7 is:
$G\approx\sqrt{\left(\frac{\Delta T_{x}}{\Delta x}\right)^{2}+\left(\frac{\Delta T_{y}}{\Delta y}\right)^{2}},$ (10)
where $\Delta x$ and $\Delta y$ are the distance between adjacent points in the smoothed temperature map in the $x$- and $y$-directions (3 $\mu$m and 3.4 $\mu$m, respectively, after smoothing the temperature map using spline interpolation in MATLAB), and $\Delta T_{x}$ and $\Delta T_{y}$ are estimated using a linear approximation of the temperature centered difference method. Similar to Equation 9, $\Delta T_{y}$ is calculated as follows:
$\Delta T_{y}=\frac{\mid T_{y}-T_{y-1}\mid+\mid T_{y}-T_{y+1}\mid}{2},$ (11)
In Equation 11, $y$ is the corresponding pixel location in the $y$-direction of the image, and $T_{y+1}$ and $T_{y-1}$ are the temperatures of adjacent pixels. Using these linear approximations, the gradient ($G$) is calculated and discussed in the main manuscript.
8. $y$-direction solidification velocity
Figure S7: Solidification velocity based on the melt pool shape. (a) illustrates the measurement of the solidification velocity ($R$) from the outermost edge of the melt pool (indicated by the red star) to the center. (b) The solidification velocity ($R$) is measured over time.
The solidification velocity at the edges of the melt pool (along the $y$-direction) moving inward can also be calculated by tracking the melt pool boundary from the side edge to the center, perpendicular to the laser scanning direction. As the melt pool moves in the laser scanning direction in each time frame, the solidification velocity, $R$, from the outermost edge of the melt pool to its center is determined. Due to the rapid movement of the melt pool, the frame rate of our experiment does not provide sufficient solidification data measurements. To address this issue, we interpolated two (for 200 mm/s scan speed), three (for 400 mm/s scan speed), and ten (for 600 mm/s scan speed) imaginary melt pool boundaries between adjacent time intervals of approximately 66 $\mu$s (with 15,136 fps for our experiment). This interpolation allows for more precise solidification rate measurements with time steps of approximately 6-20 $\mu$s in each frame. The solidification velocity was calculated as follows: At $t=t_{\textbf{0}}$, using the highest top position of the melt pool as a reference, the velocity from this maximum height at the upper edge to the melt pool center during the interval $t_{\textbf{0}}$ and $t_{\textbf{1}}$ is measured as $R_{\textbf{0-1}}$. Similarly, the solidification velocity between $t_{\textbf{1}}$ and $t_{\textbf{2}}$, is calculated as $R_{\textbf{1-2}}$.
The $y$-direction solidification velocity as a function of time is presented in Figure S7(b). It can be seen that the solidification velocity accelerates as the edge moves towards the center. The time it takes for the edges to meet at the center (the last data point in Figure S7(b)) reduces as the laser scanning speed increases. The slower laser scan speed (black data) exhibits a slower solidification velocity compared to the faster scan speed (orange and blue data). In addition, higher laser power (square-shaped data) also results in a slower solidification rate compared to the lower laser power (circle-shaped data). This observation aligns with our expectation since slower scan speeds and higher laser power retain more heat, thereby prolonging the solidification time of the melt pool. Rapid increases in solidification velocity are observed at the tail end of solidification, particularly for melt tracks with lower powers.